# Towards AI as Colleagues: Multi-Agent System Improves Structured Professional Ideation

### v1

@@ 5.3. RQ2: Impact of Multiple AI Perspectives on Creative Outcomes
Outcome Quality & Novelty Q5, Q7 5.95 ± 0.92 4.97 ± 1.16 17.0 ¡.01** 0.69
Table 6. Statistical Comparison for Performance & Integration Metrics between MultiColleagues (MC) and Baseline. Results show significantly higher ratings for MC on Creative Exploration and Outcome Quality & Novelty.
Figure 5. This figure compares discussion patterns between MultiColleagues and Baseline. Baseline generated main topics (0.78 vs. 0.36 per minute) and sub-topics (3.09 vs. 1.25 per minute) at a faster rate, but branching ratios were comparable. MultiColleagues invested significantly more time per topic (2.93 vs. 1.39 minutes) and per sub-topic (0.91 vs. 0.36 minutes), supporting a slower, more deliberate style that enabled sustained idea elaboration.
@@
Survey measures confirmed that participants experienced greater creative exploration with MultiColleagues (M = 6.00, SD = 0.87) than with the Baseline condition (M = 4.95, SD = 1.55; W = 31.5, $p$ = .018). To understand this difference, we analyzed conversational topical patterns, focusing on the pace at which new topics were introduced and the time spent developing them. In Figure 5, we found that Baseline produced new content at a faster pace, with more main topics per minute (M = 0.78 vs. 0.36, SD = 0.19 vs 0.07; W = 0.0, $p$ ¡ .001) and more sub-topics per minute (M = 3.09 vs. 1.25, SD = 0.89 vs 0.29; W = 0.0, $p$ ¡ .001). However, the branching ratio showed only a modest difference, with Baseline producing a slightly higher average branching ratio (M = 3.86, SD = 0.49) compared to MultiColleagues (M = 3.72, SD = 0.60; W = 83.0, $p$ = 0.43), which indicates both systems supported relatively similar topical divergence levels. However, for investigating participants’ time investment per topic, we found that participants in MultiColleagues resulted in more sustained exploration of each idea, where they spent over twice as much time on a single main topic (M = 2.93 vs. 1.39 min, SD = 0.71 vs. 0.36 min; W = 0.0, $p<$ .001) and on each sub-topic (M = 0.91 vs. 0.36 min, SD = 0.40 vs. 0.14 min; W = 0.0, $p<$ .001). This pattern indicates that MultiColleagues facilitated sustained exploration rather than rapid topic turnover. Interview reflections reinforced this sustained style of exploration. Several participants noted that the Baseline condition’s “large initial response” often felt like “a presentation you just read” (P19), which discouraged further exploration. By contrast, MultiColleagues encouraged participants to “proactively join to think and discuss” when only a few points were offered (P3). These results demonstrate that MultiColleagues broadened creative exploration by sustaining user-driven idea development beyond Baseline’s faster topical output.
@@
5.3.3. Conversational Rhythm Supports Idea Quality and Novelty
Self-reported measures indicated that participants generated higher-quality and more novel ideas with MultiColleagues, with ratings significantly higher in the MultiColleagues condition (M = 5.95, SD = 0.92) than in the Baseline condition (M = 4.97, SD = 1.16; W = 17.0, $p$ ¡ .01). To better assess the uniqueness of participants’ ideation, we performed TTCT analysis of originality (Guzik et al., 2023; Chi, 2024; Hadas and Hershkovitz, 2024; Fukumura and Ito, 2025) (see details in Appendix B.2). The results showed that MultiColleagues (M = 3.78, SD = 0.38) scored higher than Baseline (M = 3.59, SD = 0.44; W = 70.0, $p$ = .202), though this difference was not statistically significant. Despite this, participants consistently perceived MultiColleagues’ outputs as more original. Interview accounts explained this perception by pointing to the rhythm of idea generation. Participants emphasized that MultiColleagues generated ideas progressively, in a rhythm they described as “digestible… aligned with the rhythm of human discussion” (P19). This stepwise unfolding supported a process of guided discovery, where ideas “emerge through guided conversation” (P2) and developed like a “chain of thought” (P16). In contrast, Baseline often produced an immediate burst of ideas. Participants acknowledged these outputs as “very creative right from the start” (P2), yet also noted that the density could feel overwhelming. Several reported that this rapid surge made it harder to refine and act on ideas, compared to the more deliberate, incremental approach offered by MultiColleagues (P3, P5, P15).
@@ 6.1. Summary of Results
Our study demonstrates how MultiColleagues, a multi-agent conversational system, reshapes collaborative ideation compared to single-agent baselines. First, for RQ1 - Collaborative Experience, MultiColleagues fostered a distributed, team-like atmosphere. Participants reported stronger team-like feelings and complementary strengths across roles. Engagement also increased, with nearly twice as many utterances, longer sessions, and more words produced overall, reflecting a shift from passive receipt to facilitative coordination. For RQ2 - Creative Outcomes, multiple AI perspectives supported broader and deeper exploration: participants spent more time developing each idea topic and producing outputs judged higher in quality and novelty. For RQ3 - System Design and User Agency, MultiColleagues enabled stronger perceived control and more adaptive thinking mode support for switching between divergent and convergent thinking. Although autonomy sometimes led to conversational drift, participants valued the ability to steer discussions, regulate rhythm, and align outputs with evolving goals. Our results indicate that multi-agent systems shift ideation from tool use toward dynamics that resemble collegial teamwork. Building on these findings, we outline key design implications that translate observed collaboration patterns into actionable directions for future system design, detailed in the following sections (Figure 6).
@@
6.2. Expanding Perspectives through Proactive Multi-Agent Collaboration
@@
Our study demonstrated how MultiColleagues broadened user perspectives by combining multiple voices with proactive contributions. Instead of passively awaiting prompts, colleagues anticipated latent needs and introduced dimensions users had not explicitly considered, thereby helping them move beyond current frames of thought. At the same time, colleagues did not simply deliver isolated or premature convergence ideas (Fukumura and Ito, 2025); they engaged in internal contrasts and comparisons, allowing alternative viewpoints to emerge and be weighed before reaching the user. Building on this, we frame the MultiColleagues system as a foundation for multi-voiced conversation, where diverse AI colleagues contribute proactively and in contrast to one another, creating exchanges that are both cognitively supportive and closer to real teamwork.
6.2.1. Design Implication 1: Supporting Proactive and Multi-voiced Colleagues.
Prior work has noted the risks of homogeneous perspectives in AI systems, while also emphasizing the potential of multi-agent conversation to counteract them by surfacing diverse viewpoints and stimulating creative exchanges (Zhang et al., 2024; Pang et al., 2025; Lu et al., 2024). Our study extends this line of inquiry by showing how participants valued proactive need recognition and autonomous idea development as mechanisms to broaden perspectives and sustain engagement. To function as more autonomous, team-like collaborators, AI colleagues should not only anticipate latent user needs but also initiate their own debates and substantiate claims with evidence. Building on prior work showing how generative agents can sustain memory and social interaction over time (Park et al., 2023), future systems should also consider how designs can scale from small groups of colleagues to larger collectives while preserving the richness of multiple perspectives (Park et al., 2024). By synthesizing diverse viewpoints internally and proactively, such systems can provide more balanced and well-reasoned insights that better leverage collective intelligence.
@@ 6.3. Orchestrating Many Voices: Designing Multi-Agent Colleagues for Seamless Collaboration
6.3.1. Design Implication 2: Structuring Conversations for Seamless Flow.
@@
Managing contributions from multiple AI colleagues requires orchestration strategies that surface diverse perspectives without overwhelming users. Prior work highlights approaches such as round-table settings, sequential workflows, phased role-play, and chat-based interface for balancing diversity and order in multi-agent systems (Pang et al., 2025; Lu et al., 2024; Dibia et al., 2024; Peng et al., 2024). Our study extends this work by showing that a conversational chat style with clear turn-taking reduced barriers to entry, preserved distributed expertise, and created a digestible rhythm that supported both engagement and coordination.
Yet questions remain about how best to orchestrate multi-agent conversations across different user groups and task contexts. While sequential turn-taking fostered clarity, participants also envisioned more flexible designs, such as reconfiguring colleagues mid-discussion, introducing simultaneous responses, or staging role-based debates. Addressing the cognitive complexity of many voices further calls for visual orchestration aids, such as color-coded dialogue flows or labeled discussion maps, that provide overviews and allow lightweight steering of complex exchanges. Future research should explore how adaptive orchestration mechanisms, ranging from sequential turns to parallel exchanges, can dynamically adjust based on user preferences, task phases, or cognitive load, moving toward seamless collaboration experiences that scale across diverse ideation tasks.
6.3.2. Design Implication 3: Designing Visualization and Identity for Recognizable Colleagues
The appearance and identity cues of AI colleagues play a critical role in shaping how users perceive and engage with them, often creating a heightened sense of “being in a team” (P5). To evoke this social presence in our system, we assigned distinct profile avatars to each colleague, closely resembling profile pictures in real-world office platforms. This design choice assisted participants with a familiar visual anchor and reinforced the impression that they were collaborating with recognizable teammates rather than interacting with abstract system outputs. Looking forward, effective collaboration requires moving beyond simple profile images toward a more comprehensive identity package. Recent HCI research shows that projecting digital colleagues into office spaces (Osborne et al., 2025), rendering two-dimensional images as three-dimensional avatars (Hilpert et al., 2024), and employing AR-based or customized avatars in meetings (Sinlapanuntakul and Zachry, 2025) can strengthen social presence and improve perceptions of credibility in group work (Shamekhi et al., 2018; Seaborn, 2025). At the same time, visual identity design must avoid reinforcing stereotypes, since default gendering or normative depictions risk perpetuating bias (Aldahoul et al., 2024). Providing gender-neutral options, diverse visual styles, and customizable pronouns ensures AI colleagues are both recognizable and inclusive, which in turn supports more authentic and equitable collaboration.
@@ 6.3.2. Design Implication 3: Designing Visualization and Identity for Recognizable Colleagues
6.3.3. Design Implication 4: Orchestrating Conversational Style for Socialized and Contextual Interactions
@@
The communicative style of AI colleagues strongly influenced how smoothly information was exchanged and understood. Users’ interactions felt closer to peer-to-peer collaboration rather than mechanical delivery when AI responses adopted a more approachable and adaptable tone that could be polite, occasionally humorous, or appropriately formal. Subtle stylistic shifts created a sense of social presence across conversations that resemble real workplace exchanges and making shared content easier to follow. In addition, response length also shaped the rhythm of collaboration. Participants reflected that longer outputs were described as valuable in “focus mode” to offer more in-depth reasoning, while shorter and contextualized statements helped anchor contributions without overwhelming the discussion. Future works could focus on designing adaptive mechanisms that vary in verbosity across phases of a task, from elaboration at the outset to concise handovers later, that can support both breadth–depth exploration and efficient knowledge sharing.
6.4. From “Many Agents” to “Colleagues”: Establishing Peer-Like Roles and Trust

### v2

@@ 5.2. RQ1: Collaborative Experience with Multi-AI Colleagues
Perceived Quality & Novelty Q5, Q7 5.95 ± 0.92 4.97 ± 1.16 17.0 ¡.01** 0.69
5.2.4. Engagement and Flow Enhance Collaborative Immersion
Participants reported experiencing stronger engagement and conversational flow in MultiColleagues, with significantly higher ratings in the MultiColleagues condition (M = 5.70, SD = 1.38) compared to the Baseline condition (M = 4.45, SD = 1.57; W = 30.0, $p$ = .014). Behavioral interaction measures further reinforced this pattern (Table 5). Participants contributed nearly twice as many utterances in MultiColleagues (M = 8.35, SD = 5.79) than in the Baseline condition (M = 4.10, SD = 2.45; W = 9.5, $p$ = .001) and produced substantially more words overall (W = 18, $p$ ¡ .001), suggesting an increased engagement in ideation in the MultiColleague condition compared to the Baseline. Sessions also lasted longer in MultiColleagues (M = 12.90, SD = 2.80) compared to Baseline (M = 9.80, SD = 2.30; W = 33, $p$ = .006), suggesting that participants remained more cognitively and temporally invested when coordinating multiple voices. Interview reflections echoed these dynamics. P5 explained, “I feel like in my discussions, I’m constantly prompting new information and asking for new information, kind of wanting to lead the discussion. I feel I’m more like a facilitator.” Others emphasized how the incremental delivery kept them attentive and invested (P3, P10, P12). While MultiColleagues fostered ongoing involvement, Baseline condition’s one-shot responses were often described as efficient but less engaging. P19 described the interaction as “relatively one-way, like a presentation… I just pick what I want,” and P8 admitting, “I didn’t feel like we were collaborating… I just wanted to hear its perspective.” These findings show that MultiColleagues’ structured rhythm sustained deeper immersion and engagement, while Baseline delivered efficiency at the cost of shallower participation.
5.3. RQ2: Perceived Impact on Ideation Process and Outcome Quality
To address RQ2, we examined participants’ evaluations of creative outcomes along three dimensions: creative exploration, process enrichment, and perceived outcome quality and novelty. Participants rated MultiColleagues more favorably than Baseline across all three dimensions, with significant differences for creative exploration and perceived outcome quality and novelty (Table 6).
Figure 5. This figure compares users’ idea development patterns across MultiColleagues (MC) and Baseline (BL). MC produced a higher topic branching (6.45 vs. 3.50, $p<0.001$) and higher concepts production rate (2.05 vs. 1.26 concepts per minute, $p<0.05)$.
Statistical comparison of users’ ideation development between Baseline and MultiColleagues. The figure shows that MultiColleagues demonstrate significantly more topic branching and concept production, which supports broader and more in-depth ideation.
@@ 5.3. RQ2: Perceived Impact on Ideation Process and Outcome Quality
Survey measures confirmed that participants experienced greater creative exploration with MultiColleagues (M = 6.00, SD = 0.87) than with the Baseline condition (M = 4.95, SD = 1.55; W = 31.5, $p$ = .018). Further, statistical tests did not reveal significant associations between number of personas selected and creativity scores ($r=0.223,p=0.345$) or occupational background ($H=1.292,p=0.524$). These findings suggest that inclusion of a good enough number of different perspectives would facilitate creativity; however, increasing the number of personas solely did not guarantee higher creativity. To further examine this difference, we analyzed the topical structure and conceptual content of user contributions. Topic branching captured how many parallel idea directions participants pursued, while concepts production reflected the amount of concepts (i.e., nouns, (Kintsch, 1998).) people produced in their language discourse. Figure 5 shows that MultiColleagues supported broader and more in-depth ideation than the baseline. Participants generated significantly higher topic branching with MultiColleagues (M = 6.45, SD = 2.54) than Baseline (M = 3.50, SD = 1.79, $t(19)=5.15,p<.001$), indicating that they explored more parallel lines of ideas in terms of broader elaboration of the topics. Participants also generated more concepts in their discussion, suggesting more in-depth reflections, as shown in a higher concept production rate (M = 2.05 concepts/min, SD = 0.81) in MultiColleagues than in Baseline (M = 1.26, SD = 0.67, $t(19)=2.59,p<.05$). These results show MultiColleagues encouraged both greater structural divergence and richer conceptual expression during ideation. Interview reflections reinforced this pattern. Several participants noted that the Baseline condition’s “large initial response” often felt like “a presentation you just read” (P19), which reduced their motivation to extend the ideas in breadth or depth. In contrast, MultiColleagues encouraged participants to “proactively join to think and discuss” when only a few points were offered (P3). These results demonstrate that MultiColleagues broadened creative exploration by enabling participants to pursue more directions and engage in more conceptually rich discourse.
@@
5.3.3. Conversational Rhythm Shapes Perceived Idea Quality and Novelty
Self-reported measures indicated that participants perceived their ideas as higher-quality and more novel with MultiColleagues, with ratings significantly higher in the MultiColleagues condition (M = 5.95, SD = 0.92) than in the Baseline condition (M = 4.97, SD = 1.16; W = 17.0, $p$ ¡ .01).
To complement these perception-based measures, we additionally examined idea originality using a TTCT-inspired scoring approach (Guzik et al., 2023; Chi, 2024; Hadas and Hershkovitz, 2024; Fukumura and Ito, 2025). TTCT evaluates creativity across fluency, flexibility, originality, and elaboration. In this study, fluency and flexibility were not included, as previous works demonstrated that LLMs tend to produce a high volume of responses that artificially inflate fluency scores, while flexibility is strongly confounded by fluency and thus offers limited validity as an independent measure (Guzik et al., 2023; Chi, 2024; Hadas and Hershkovitz, 2024). Instead, we focused on originality as the core dimension. Following prior work (Fukumura and Ito, 2025), originality ratings were generated using GPT-5 with the same validated evaluation prompt applied to participants’ full conversations. Each conversation was evaluated on a standardized 5-point rubric, with three independent runs per idea set, and the averaged score was used for originality analysis. The results showed a trend that MultiColleagues (M = 3.78, SD = 0.38) scored higher than Baseline (M = 3.59, SD = 0.44; W = 70.0, $p$ = .202), but this difference was not statistically significant, diverging from the perceived novelty reported by participants. Interview accounts help explain this divergence by highlighting differences in the rhythm of idea generation. Participants emphasized that MultiColleagues generated ideas progressively, in a rhythm they described as “digestible… aligned with the rhythm of human discussion” (P19). This stepwise unfolding supported a process of guided discovery, where ideas “emerge through guided conversation” (P2) and developed like a “chain of thought” (P16). In contrast, Baseline often produced an immediate burst of ideas. Participants acknowledged these outputs as “very creative right from the start” (P2), yet also noted that density could feel overwhelming. Several reported this rapid surge made it harder to refine and act on ideas, compared with MultiColleagues’ more deliberate, step-by-step approach (P3, P5, P15).
@@ 5.4.3. Adaptive Thinking Mode Strengthens Flexibility in Ideation
Design implications for human–multi-agent collaboration. The figure presents 5 implications that mark the shift from AI as task-oriented tools to AI as proactive, accountable, and socially embedded colleagues, highlighting considerations in agents’ proactivity, persona dimensions, orchestrations, conversational style, and reliability.
@@ 6.1. Summary of Results
Our study demonstrates how MultiColleagues, a multi-agent conversational system, reshapes collaborative ideation processes compared to a single-agent baseline. First, for RQ1 - Collaborative Experience, MultiColleagues fostered a distributed, team-like atmosphere. Participants reported stronger team-like feelings and complementary strengths across roles. Engagement also increased, with nearly twice as many utterances, longer sessions, and more words produced overall, reflecting a shift from passive receipt to facilitative coordination. For RQ2 - Perceived Creative Outcomes, the exposure to multiple AI colleague perspectives supported broader and deeper exploration, participants spent more time developing each idea topic and reported higher perceived outcome quality and novelty. For RQ3 - System Design and User Agency, MultiColleagues enabled stronger perceived control and more adaptive thinking mode support for switching between divergent and convergent thinking. Although autonomy sometimes led to conversational drift, participants valued the ability to steer discussions, regulate rhythm, and align outputs with evolving goals. Our results indicate that multi-agent systems shift ideation from tool use toward dynamics that resemble collegial teamwork. Building on these findings, we outline key design implications that translate observed collaboration patterns into actionable directions for future system design, detailed in the following sections (Figure 6).
@@
6.2. Expanding Comprehensive Perspectives through Proactive Multi-Agent Collaboration
@@
Our study demonstrated how MultiColleagues broadened user perspectives by combining multiple voices with proactive contributions. Instead of passively awaiting prompts, colleagues anticipated latent needs and introduced dimensions users had not explicitly considered, thereby helping them move beyond current frames of thought. Building on this, we frame MultiColleagues system as a foundation for multi-voiced conversation, where diverse AI colleagues contribute proactively and in contrast to one another, creating exchanges that are both cognitively supportive and closer to real teamwork.
6.2.1. Design Implication 1: Proactive Colleagues Open New Perspectives
Prior work has noted the risks of homogeneous perspectives in AI systems, while also emphasizing the potential of multi-agent conversation to counteract them by surfacing diverse viewpoints and stimulating creative exchanges (Zhang et al., 2024; Pang et al., 2025; Lu et al., 2024). Our findings extend this work by showing that users valued not only the presence of multiple agents but also their proactive behavior. Colleagues that raised overlooked considerations or challenged premature convergence helped broaden perspectives and sustain engagement (Zhang et al., 2021). In our prototype, each colleague followed role based instructions to monitor the discussion and intervene when they could add a new angle or critique. Colleagues also contrasted and compared one another’s ideas, which allowed alternative viewpoints to surface and be considered before reaching the user. We also designed the facilitator who served as the main conversational lead, tracking how the conversation unfolded and indicating when the group should continue exploring or shift toward refinement.
Future systems could introduce more proactive colleagues that recognize when and how to take initiative in the discussion. Colleagues and facilitators could label the function of their contributions, for instance by marking a message as opening a new issue or responding with a challenge to a prior idea. By shaping when colleagues and facilitators intervene, and how their perspectives are contrasted and summarized, designers can better harness proactive multi-agent collaboration as a resource for broadening perspectives rather than overwhelming users.
6.2.2. Design Implication 2: From Occupational Role-play to Domain Expert Teamwork
In this study, personas were primarily differentiated by occupation, which provided distinct perspectives but offered limited guidance on how colleagues should contribute across exploration and evaluation stages. Interviews indicated that this occupation focused framing sometimes kept colleagues within narrow professional boundaries and left participants uncertain about who had sufficient expertise or authority for specific issues (P2, P7, P18). P7 explicitly noted that occupational breadth was not always the most relevant form of diversity. When his discussion focused on a specific machine learning subproblem, bringing in more cross occupational colleagues did not help his ideation. Instead, he emphasized the value of differentiating colleagues within the same domain by specialization focus and level of technical authority so that some colleagues could provide rigorous critique and others could take responsibility for synthesis during convergence.
Recent multi-agent ideation work also finds that increasing the number of agents is insufficient. Idea quality depends on agents with adequate expertise who can sustain critique and refinement over time (Chen et al., 2025a). For system design, this implies encoding clearer differences in seniority and decision rights among colleagues, so evaluation feedback is appropriately confident and responsibility for synthesis is unambiguous. Work on conversational personas and framing shows that persona cues and AI’s presented role can shape user expectations and interaction strategies, therefore should be designed to avoid deceptive signaling or reinforcing stereotypes (Pradhan and Lazar, 2021; Liao and Sundar, 2021). Future systems could treat colleague configurations as explicit, adjustable dimensions and examine how they shape teammate selection, for example preferring a smaller set of high authority evaluators for convergence and adding more contrasting colleagues for early stage exploration (Chen et al., 2025a).
@@ 6.3. Orchestrating Many Voices: Designing Multi-Agent Colleagues for Seamless Collaboration
6.3.1. Design Implication 3: Initiating Thread Tracking and Adaptive Turn Taking in Multi-Agent Chat.
@@
Managing contributions from multiple AI colleagues requires orchestration strategies that surface diverse perspectives without overwhelming users. Prior work highlights approaches such as round-table settings, sequential workflows, phased role-play, and chat-based interface for balancing diversity and order in multi-agent systems (Pang et al., 2025; Lu et al., 2024; Dibia et al., 2024; Peng et al., 2025). Our study extends this line of work by showing that a conversational chat style with clear turn taking lowered barriers to entry and preserved a distributed sense of expertise, which in turn produced a digestible rhythm that supported engagement and coordination during short, time bounded ideation sessions.
As the number of AI colleagues grows, legibility and fast attribution become central. Users need to “know which one to talk to” (P3), follow the development of each thread, and return to specific contributions when ideas converge. In our prototype, distinct avatars and role labels prompted participants to treat colleagues as recognizable teammates rather than a single undifferentiated system voice, which aligns with research that links visual embodiment to stronger social presence and coordination in group work (Shamekhi et al., 2018; Hilpert et al., 2024). Building on this, multi-agent interfaces could pair identity cues with lightweight visualization, such as color coded dialogue flows or labeled discussion maps, thus users can trace how proposals and critiques evolve. Participants also imagined more flexible and controllable interaction patterns. They described needs such as reconfiguring colleagues mid discussion (P1, P5), inviting simultaneous responses for comparison (P2), and setting up focused debates between evaluators and proponents (P4, P7). These expectations point toward identity based steering hooks where users can click a colleague to request targeted critique, re-invite a silent expert, mute a dominant voice, or change who speaks first when the team moves from exploration to convergence.
From the system side, our prototype used prompt level instructions to coordinate turns. Each colleague received the shared conversational history and was asked to decide whether to speak in the current round. The orchestrator then ranks the colleagues with higher reported priority for each turn that produces a manageable flow. Beyond this basic scheme, adaptive orchestration can draw on a richer set of conversational signals. Prior works show that interaction traces and user cues can guide when agents take or yield the floor (Hu et al., 2025; Xu et al., 2013). Building on this, multi-agent systems can monitor which roles participate and how often users seek clarification, then adapt the mix of colleagues and the pacing of turns to fit the current phase of work. Future work could formalize such adjustments as phase sensitive orchestration policies, extending existing models of turn taking and turn detection in multi-party systems. (Aldahoul et al., 2024).
@@
6.3.2. Design Implication 4: Adaptive Conversational Style and Response Length Facilitate Socialized Interactions
@@
The communicative style of AI colleagues strongly influenced how smoothly information was exchanged and understood. Users’ interactions felt closer to peer-to-peer collaboration rather than mechanical delivery when AI responses adopted a more approachable and adaptable tone that could be polite, occasionally humorous, or appropriately formal. Subtle stylistic shifts created a sense of social presence across conversations that resemble real workplace exchanges and making shared content easier to follow. In addition, response length also shaped the rhythm of collaboration. Participants reflected that longer outputs were described as valuable in “focus mode” to offer more in-depth reasoning, while shorter and contextualized statements helped anchor contributions without overwhelming the discussion (P20). Future works could focus on designing adaptive mechanisms that vary in verbosity across phases of a task, from elaboration at the outset to concise handovers later, that can support both breadth–depth exploration and efficient knowledge sharing.
6.4. From “Many Agents” to Colleagues: Role Stability and Reliable Use